The output voltage of an AC generator is given by ?v = (120 V) sin (50 ?t). The generator is connected across a 25.0-? resistor.

(a) By inspection, what is the maximum voltage?
V

(b) By inspection, what is the frequency?
Hz

(c) Find the rms voltage across the resistor.
V

(d) Find the rms current in the resistor.
A

(e) Find the maximum current in the resistor.
A

(f) Find the power delivered to the resistor.
W

(g) Find the current when t = 0.0050 s.
A

(h) Should the argument of the sine function be in degrees or radians?
The argument of the sine function should be in radians.The argument of the sine function should be in degrees.
Answer #1

a) maximum voltage= 120V


b) frequency= 2?*50?=314.15Hz


c)rms voltage across the resistor=120/2^1/2=84.85V


d)rms current in the resistor=84.85/25=3.39A


e) Maximum Current in the resistor=120/25=4.8A


f) power delivered to the resistor= 3.39^2 * 25=287.3W


g) when t = 0.0050 s, Current= (120 V) sin (50?*005)/25 = 0.0657A


h) the arguement of sine function nshould be in radian